OpenStudy (anonymous):

Find the smallest possible value of \( x^2 + bx + c \) and of \( ax^2 + bx + c \) for a > 0 I've shown by completing the square that \(x^2 + bx + c = (x + b/2)^2 + c - b^2 / 4 > 0\)

OpenStudy (anonymous):

stuck :-D

OpenStudy (anonymous):

for \(ax^2 + bx + c = a(x^2 + b/a x + c/a) = a((x + b/2a)^2 - b^2/4a + c > 0)\)

OpenStudy (anonymous):

(x+b/2)2 is either 0 or positive.

OpenStudy (anonymous):

can't you just find vertex?

OpenStudy (anonymous):

yea that works

OpenStudy (anonymous):

what's the vertex denoted by the above two equations (after I completed their square)

OpenStudy (anonymous):

*facepalm seems to be c−b2/4 for the first one, and −b2/4a+c for the second one